# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

### v2

@@ 4. Results
We now present quantitative results organized by the research questions that we aim to answer. We report results by answering the five research questions ,across the settings mentioned above, based on our experiments backed by tables and figures.
RQ.1. Does Role Identity Bias Prevent Nash Equilibrium Achievement in Economic Scenarios?
Answer: Yes. When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning. Fig. 3 shows our findings. When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information. Models select Green Transition actions (20-65%, Fig. 2 shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
Table 2. Environmental Scenario Nash Equilibrium Rates: With Personas vs. Without Personas. This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
Model Payoffs With Persona Without Persona Difference
Llama-8B Hidden 100.0% (205/205) 97.6% (200/205) -2.4 pp
Llama-8B Visible 99.5% (204/205) 93.2% (191/205) -6.3 pp
Mistral-7B Hidden 100.0% (205/205) 98.0% (201/205) -2.0 pp
Qwen 32B Hidden CoT 89.3% (183/205) 96.1% (197/205) +6.8 pp
Qwen 7B Hidden CoT 95.6% (196/205) 95.6% (196/205) 0.0 pp
Table 3. Economic Scenario Nash Equilibrium Rates: Hidden vs Visible Payoffs. This table compares Nash equilibrium rates in economic scenarios (where Tragedy is payoff-optimal) between Hidden and Visible payoff conditions. Each row shows a different model-persona combination.
Model Persona Hidden Visible Difference
Llama-8B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Llama-8B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
Mistral-7B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Mistral-7B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
Qwen-32B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Qwen-32B No 0.0% (0/60) 90.0% (54/60) +90.0 pp
Qwen-7B Yes 6.7% (4/60) 0.0% (0/60) -6.7 pp
Qwen-7B No 6.7% (4/60) 65.0% (39/60) +58.3 pp
Our findings show that role identity bias prevents Nash equilibrium when personas are present, causing models to align actions with persona expectations rather than payoff-optimal Tragedy actions. Chain-of-thought analysis reveals that with personas, models show more “short-term” and “social-moral” reasoning patterns, indicating identity-driven rather than payoff-optimal decision-making. CoT analysis also reveals different reasoning patterns between models. Llama and Mistral show consistent patterns regardless of payoff visibility, while Qwen models show some payoff-focused keywords but still cannot overcome identity bias to select Tragedy equilibrium.
RQ.2. How does removing personas affect Nash equilibrium achievement?
Answer: Removing personas enables partial Tragedy selection in economic scenarios but reduces overall Nash rates by 1.9-7.3 percentage points. Table 1 compares overall Nash equilibrium rates across all scenarios (economic and environmental) between conditions with and without personas. This table demonstrates that removing personas generally reduces Nash rates for most models, except Qwen-32B with hidden payoffs, which shows a slight increase.
Personas both enable and constrain Nash equilibrium achievement. Overall Nash rates decrease by 1.9-4.9 pp for Llama and by 1.6 pp for Mistral when personas are removed, and environmental scenario Nash rates decrease by 2-6 pp (Table 2) without personas for most models, indicating personas enable better coordination in Green Transition scenarios. CoT analysis shows that with personas, models exhibit more “social-moral” reasoning, which aligns with socially preferred Green Transition outcomes. However, Qwen-32B shows the opposite pattern, where Nash rates increase by 5.2 pp (Table 1) without personas, suggesting personas constrain Qwen’s ability to reach Nash equilibrium. CoT analysis reveals that removing personas shifts reasoning from “social-moral” and “short-term” patterns to “strategic” and “long-term” thinking for Qwen models, indicating a shift from identity-driven to payoff-optimal reasoning.
In economic scenarios, removing personas reduces Green Transition selection by 10-37 percentage points (see Table 10 in the Appendix), indicating partial Tragedy selection when personas are absent. However, removing personas alone is insufficient. Without personas but with hidden payoffs, all models still achieve 0% Nash in economic scenarios. Only Qwen models with visible payoffs and no personas can select Tragedy equilibrium (65-90% Nash, Fig. 3), demonstrating that both removing personas and providing explicit payoffs are necessary for payoff-optimal reasoning.
RQ.3. Does information presentation format (hidden vs visible payoffs) affect Nash equilibrium achievement?
Answer: Yes. Hidden payoffs achieve similar or higher Nash rates than visible payoffs when personas are present, but visible payoffs enable Tragedy equilibrium selection when personas are removed (for Qwen models).
Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs. The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%. Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern. The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings. This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model architectures and experimental conditions.
Figure 2. Green Transition Action Selection Rates in Economic Scenarios. This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where the Tragedy of Commons is the payoff-optimal equilibrium.
Refer to caption Heatmap showing Nash equilibrium rates across four models and four condition combinations A heatmap displaying Nash equilibrium rates (0-100%) for economic scenarios across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) and four experimental conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona. The heatmap uses a blue color gradient where lighter shades represent 0% Nash rates and darker shades represent higher rates. Most cells show 0.0% (light blue), with two notable exceptions: Qwen 32B in the Visible Payoff without Persona condition shows 90.0% (dark blue), and Qwen 7B in the same condition shows 65.0% (medium-dark blue). This visualization clearly demonstrates that only Qwen models achieve high Nash equilibrium rates in economic scenarios, and only when both conditions are met: payoffs are visible and personas are removed. All other model-condition combinations result in 0% Nash rates, with Qwen 7B showing minor exceptions (6.7%) in Hidden Payoff conditions regardless of persona presence.
Figure 3. Economic Scenarios Nash Equilibrium Rates by Model and Condition. This heatmap visualizes Nash equilibrium achievement rates in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium across all four models and four experimental conditions. The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
Table 4. Equilibrium Selection Patterns: With Personas Condition. Note: “Green % of Nash” shows the percentage of Nash equilibria that are Green Transition. The denominators vary because different models/variants achieve Nash equilibrium at different rates (see “Overall Nash” column), but when Nash is achieved, it is always Green Transition (100%).
Model Variant Overall Nash Green Transition Rate Tragedy Rate Green % of Nash
Llama-8B Hidden 77.4% (205/265) 87.5% (232/265) 0.0% (0/265) 100.0% (205/205)
Llama-8B Visible 77.0% (204/265) 77.0% (204/265) 0.0% (0/265) 100.0% (204/204)
Mistral-7B Hidden 77.4% (205/265) 90.6% (240/265) 0.0% (0/265) 100.0% (205/205)
Mistral-7B Visible 69.8% (185/265) 69.8% (185/265) 0.0% (0/265) 100.0% (185/185)
Qwen-32B Hidden 69.1% (183/265) 69.8% (185/265) 0.0% (0/265) 100.0% (183/183)
Qwen-32B Visible 57.7% (153/265) 57.7% (153/265) 0.0% (0/265) 100.0% (153/153)
Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
Qwen-7B Visible 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions. When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models. This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices. CoT analysis confirms this claim. With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions.
This pattern reverses when personas are removed. Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference). CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization. This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed. However, this capability is model-dependent. Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
The effect of visible payoffs also differs across scenario types. Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models, dropping from 96% to 17% for Qwen 32B, suggesting a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios. Mistral shows the opposite pattern, with visible payoffs and no personas, reducing environmental Nash from 98% to 50% (Table 5), while economic Nash remains at 0% (Table 3), suggesting Mistral cannot leverage explicit payoffs effectively regardless of scenario type.
Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs. Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help.
Model Persona Hidden Visible Difference
Llama-8B Yes 100.0% (205/205) 99.5% (204/205) -0.5 pp
Llama-8B No 97.6% (200/205) 93.2% (191/205) -4.4 pp
Mistral-7B Yes 100.0% (205/205) 90.2% (185/205) -9.8 pp
Mistral-7B No 98.0% (201/205) 50.2% (103/205) -47.8 pp
Qwen 32B Yes 89.3% (183/205) 74.6% (153/205) -14.7 pp
Qwen 32B No 96.1% (197/205) 16.6% (34/205) -79.5 pp
Qwen 7B Yes 95.6% (196/205) 97.6% (200/205) +2.0 pp
Qwen 7B No 95.6% (196/205) 52.7% (108/205) -42.9 pp
RQ.4. Are persona effects model-dependent?
Answer: Yes. Qwen models show stronger persona dependence (9.7-10.0 pp average reduction in Nash rates without personas), while Llama-8B shows minimal dependence (3.4 pp average, range 1.9-4.9 pp). [^5] See Tables 6(b)6(a), and 12 in the Appendix for detailed persona effect statistics..
Table 6(a) quantifies the magnitude of persona effects on overall Nash equilibrium rates. All models show negative persona effects in the visible payoffs variant (-5 to -31 pp), but the mechanisms differ.
For Qwen models, removing personas enables payoff-optimal Tragedy equilibrium in economic scenarios but reduces Green Transition equilibrium in environmental scenarios when visible payoffs are present (96.1% $\to$ 16.6% Nash for Qwen-32B, Table 5), resulting in net negative overall effects (-20-25 pp, Table 6(a)). For Llama/Mistral, removing personas only reduces environmental Nash equilibrium without enabling economic Nash (0% remains), reflecting pure Nash rate loss. Llama-8B shows minimal persona dependence (1.9-4.9 pp reduction), while Mistral-7B shows large effects (-30.9 pp) due to reduced environmental Nash rates.
Analyzing the Chain-Of-Thought of the models shows how differently they reason. Llama maintains consistent reasoning patterns across conditions (minimal keyword differences), suggesting its reasoning is relatively unaffected by persona presence. Mistral shows payoff optimization keywords but cannot identify Tragedy equilibrium regardless of condition. Qwen models show dramatic shifts with and without personas. With personas, they exhibit “short-term” reasoning; without personas and with visible payoffs, they shift to “strategic” (368 mentions for Qwen-32B) and “long-term” (283 mentions) reasoning. 99.7% of the rationales [^6] See Section E for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning.
Additionally, model size does not predict persona dependence. Both Qwen-7B and Qwen-32B show similar reasoning pattern shifts, but Qwen-32B achieves higher Nash rates due to stronger payoff optimization.
Refer to caption Two-panel stacked bar chart showing Green Transition and Tragedy of Commons percentages A two-panel stacked bar chart displaying the percentage of Nash equilibria that are Green Transition versus Tragedy of Commons across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) in the without personas condition. The left panel shows the Hidden Payoff condition, where all models achieve 100% Green Transition (blue bars with diagonal stripes). The right panel shows the Visible Payoff condition, where Llama 8B and Mistral 7B maintain 100% Green Transition, while Qwen 32B shows 11.4% Green Transition (green) and 88.6% Tragedy of Commons (orange-red), and Qwen 7B shows 89.8% Green Transition and 10.2% Tragedy of Commons. Each bar segment is labeled with its percentage value in white text. The legend is positioned below the figure, showing Green Transition in blue with diagonal stripes and Tragedy of Commons in orange-red with dots.
Figure 4. Equilibrium Selection Patterns: Green Transition vs Tragedy of Commons (Without Personas Condition). This figure shows the distribution of Nash equilibria between Green Transition and Tragedy of Commons outcomes across all four models in the no-personas condition.
RQ.5. Do personas enable equilibrium selection beyond pure payoff maximization?
Answer: Yes. When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased. Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants. Fig. 4 reveals model-dependent patterns when personas are removed. For Qwen-32B with visible payoffs, only 11.4% of Nash equilibria are Green Transition (10/88), meaning 88.6% are Tragedy, demonstrating complete reversal from the persona condition where 100% were Green Transition. For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 3 means most economic Nash is Tragedy. For Llama and Mistral, 100% of Nash equilibria remain Green Transition even without personas, indicating they cannot identify Tragedy equilibrium regardless of condition.
Table 6. Persona Effects on Nash Equilibrium Rates: Overall (left) and Economic Scenarios (right). Left table shows overall persona effects across all scenarios (economic and environmental combined). Right table shows persona effects specifically for economic scenarios (where Tragedy is payoff-optimal). Negative values (percentage points) indicate personas help and positive values (percentage points) indicate personas hurt.
(a) Overall Nash Equilibrium Rates and their dependence on personas for various LLM architectures.
Model Hidden CoT Visible CoT Average
Llama 8B -1.9 pp -4.9 pp -3.4 pp
Mistral 7B -1.6 pp -30.9 pp -16.3 pp
Qwen 32B +5.2 pp -24.5 pp -9.7 pp
Qwen 7B 0.0 pp -20.0 pp -10.0 pp
(b) Economic Scenarios
Model Hidden CoT Visible CoT Average
Llama 8B 0.0 pp 0.0 pp 0.0 pp
Mistral 7B 0.0 pp 0.0 pp 0.0 pp
Qwen 32B 0.0 pp +90.0 pp +45.0 pp
Qwen 7B 0.0 pp +65.0 pp +32.5 pp
CoT analysis confirms the mechanism. With personas, models show “social-moral” and “short-term” reasoning, indicating identity-driven decision-making. Without personas and with visible payoffs, Qwen models shift to nearly universal “game-theoretic” reasoning (99.7% of rationales for Qwen-32B), with “strategic” and “long-term” keywords unique to this condition, demonstrating a complete shift from identity-driven to payoff-optimal reasoning that enables equilibrium selection reversal. Llama and Mistral maintain consistent reasoning patterns regardless of condition, explaining their inability to shift equilibrium selection.
@@ 5. Discussion
We ground our discussion in specific empirical findings from above and discuss what these results imply for developers, practitioners, and system design.
5.0.1. Role Identity Bias Fundamentally Alters Strategic Reasoning
Personas prevent models from identifying payoff-optimal equilibria in economic scenarios, with 12 out of 16 experiments achieving 0% Nash equilibrium despite explicit payoff information (Fig. 3). Persona expectations override explicit incentive structures, leading to reasoning patterns shifting from strategic optimization to role-aligned decision-making. This demonstrates that personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, consistent with prior findings that role prompts and system-level personas systematically bias strategic decision-making in LLMs (Neumann et al., 2025; Herr et al., 2024).
Our results show that personas are not just cosmetic enhancements but design choices that fundamentally alter system behavior. Personas systematically bias equilibrium selection toward socially preferred outcomes (100% Green Transition with personas, Table 4). Developers should explicitly document and disclose persona choices, as they determine system behavior. Users require clear information about system behavior, particularly in applications such as policy analysis or economic simulation where they may assume agents are behaving optimally according to their payoffs.
5.0.2. Payoff-Optimal Reasoning Requires Removing Personas and Providing Explicit Payoffs

### v3

@@ 4. Results
We report results separately for Green-dominant (environmental) scenarios (41 scenarios, 205 runs per condition) and Tragedy-dominant (economic) scenarios (12 scenarios, 60 runs per condition). We first examine Nash attainment across conditions, then equilibrium selection conditional on attainment, and finally prompt sensitivity.
4.1. Persona and Payoff Effects on Nash Attainment Depend on Scenario Structure and Model Family
The same representational change can either increase or decrease Nash equilibrium attainment depending on the strategic structure of the scenario. Figure 2 shows this interaction across persona and payoff visibility conditions for both scenario types.
(a) Green-dominant Scenarios (Environmental)
(b) Tragedy-dominant Scenarios (Economic)
Figure 2. Nash equilibrium attainment across persona and payoff visibility conditions. Left: Green-dominant scenarios where coordinated pro-environmental action forms the equilibrium. Right: Tragedy-dominant scenarios where individually beneficial actions form the payoff-optimal equilibrium. Visible payoffs improve equilibrium attainment only when persona descriptions are removed, and primarily for Qwen models. Chi-square tests confirm significant differences across conditions for Qwen models in both scenario types ($p\ll 0.001$). Llama-8B and Mistral-7B show no variation in Tragedy-dominant scenarios (0% Nash across all conditions).
[…]
In Tragedy-dominant scenarios, models almost never reach Nash equilibrium unless personas are removed and payoff information is explicitly provided. Under the Visible Payoff + No Persona condition, Qwen-7B attains equilibrium in 65% of runs and Qwen-32B 90% (Qwen-32B: $\chi^{2}(3)=209.03$, $p<.001$, $V=0.93$; Qwen-7B: $\chi^{2}(3)=105.91$, $p<.001$, $V=0.66$), while all other conditions remain at or near 0%. Pairwise comparisons confirm that all significant differences are driven by contrasts involving this condition ($p<0.001$, Fisher exact tests, Holm-corrected). Llama-8B and Mistral-7B fail to reach equilibrium in any condition (0% across all cells, degenerate tests with no variation).
In Green-dominant scenarios, equilibrium rates are high across most conditions (74.6–100%), indicating that role semantics and narrative context align agents toward the Green Transition outcome without explicit payoff information. However, when personas are removed and payoffs are made visible, coordination breaks down to different degrees across models. Qwen-32B drops from 96.1% under Hidden Payoff + No Persona to 16.6% under Visible Payoff + No Persona (e.g., 96.1% to 16.6% for Qwen-32B; omnibus $\chi^{2}(3)=377.06$, $p<.001$, $V=0.68$; pairwise contrasts involving Visible++No persona: $p<.001$, Fisher exact tests, Holm-corrected). Qwen-7B drops from 95.6% to 52.7%. Mistral-7B drops from 98.0% to 50.2%. Llama-8B shows the smallest change, from 97.6% to 93.2%.
In both scenario types, the underlying pattern is the same: without persona framing, agent behavior becomes more responsive to explicit payoff structure. This improves attainment when the equilibrium requires payoff-dominant actions but disrupts coordination when role semantics already support a shared outcome. The sensitivity to these representational choices varies across model families in distinct ways. Qwen models show the largest swings in both scenario types, shifting between role-aligned and payoff-aligned behavior depending on condition. Mistral-7B shows strong sensitivity in Green-dominant scenarios (47.8 pp drop) but remains at 0% in all Tragedy-dominant conditions. Llama-8B shows minimal variation across all conditions in both scenario types.
Figure 3. Equilibrium selection among Nash outcomes for Qwen models across conditions. Under the first three conditions, nearly all Nash outcomes are Green Transition. Under Visible Payoff + No Persona, approximately 40% shift to Tragedy of the Commons, indicating payoff-aligned behavior when role-based framing is removed ($\chi^{2}=430.38$, $p<10^{-90}$; pairwise Fisher tests, Holm-corrected, confirm this condition differs from all others)
A horizontal stacked bar chart showing equilibrium selection among Nash outcomes for Qwen models across four conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona. Each bar represents the percentage of outcomes classified as Green Transition (green segment) and Tragedy of the Commons (red segment). In the first three conditions, nearly all outcomes are Green Transition, with negligible Tragedy. In the Visible Payoff without Persona condition, approximately 60 percent of outcomes are Green Transition and 40 percent are Tragedy of the Commons, showing a clear shift in equilibrium selection.
4.2. Personas Shift Equilibrium Selection Toward Socially Preferred Outcomes
Personas influence not only whether equilibrium is reached, but which equilibrium is selected. With personas, equilibrium selection is entirely skewed toward socially preferred outcomes, while removing personas enables a shift toward payoff-dominant strategies.
Across all models and conditions with personas, nearly all Nash equilibria correspond to the Green Transition (over 99%). In contrast, under Visible Payoff + No Persona, Qwen models exhibit a substantial shift: only 11.4% of equilibria remain Green for Qwen-32B, with 88.6% shifting to Tragedy of the Commons, while Qwen-7B shows a smaller but still substantial shift (60.4% Green, 39.6% Tragedy). This represents changes of up to 90 percentage points in equilibrium selection ($\chi^{2}=430.38$, $p<10^{-90}$; pairwise Fisher tests, Holm-corrected, confirm that the Visible + No Persona condition differs from all others)
Llama-8B and Mistral-7B do not exhibit this shift, maintaining 100% Green Transition outcomes even without personas. This is consistent with the attainment findings above: these models never attain the Tragedy equilibrium in any condition, so their selection is restricted to Green by default.
Figure 3 shows this pattern for the Qwen models. Under the first three conditions, nearly all outcomes are Green Transition. Under Visible Payoff + No Persona, approximately 40% of outcomes shift to Tragedy of the Commons, indicating that agents move toward payoff-aligned strategies at the cost of coordinated outcomes.
4.3. Persona Wording Induces Systematic Variability in Outcomes
Even when incentives and game structure are fixed, varying persona descriptions leads to different strategic outcomes, indicating that prompt formulation alone can alter equilibrium behavior.
Figure 4 shows that scenarios produce between 2.2 and 2.4 distinct outcomes on average across models, with maximum values reaching up to 5 outcomes in some cases. This indicates that different persona variants can lead to multiple distinct equilibria within the same scenario.
This variability is not driven by stochastic differences across models. A Kruskal-Wallis test shows no significant difference in variability across model families ($p=0.27$), indicating that all models exhibit similar levels of prompt sensitivity. Instead, the observed variation reflects systematic shifts induced by persona framing. Even when the underlying payoff structure is unchanged, different persona descriptions can lead agents to select different action profiles and equilibria. This reinforces the finding from the main experimental conditions that representational choices meaningfully influence behavior in multi-agent LLM systems.
Figure 4. Prompt sensitivity across persona variants. Bars show the average number of distinct outcome classifications per scenario across seven persona prompt variants. Error bars show the minimum and maximum across scenarios. Scenarios typically produce more than two distinct outcomes on average, with some reaching up to five.
A horizontal bar chart showing prompt sensitivity across four models: Llama-8B, Mistral-7B, Qwen-32B, and Qwen-7B. The x-axis represents the average number of distinct outcome classifications per scenario, and the y-axis lists the models. Each bar shows the average value, with labels around 2.2 to 2.4 across models. Error bars extend from the minimum to the maximum number of distinct outcomes observed per scenario. For all models, the minimum is around 1 distinct outcome, while the maximum ranges from about 3 to 5, with Qwen-7B showing the widest range. This indicates that different persona prompt variants can lead to multiple distinct outcomes within the same scenario.
@@ 5. Discussion
We ground our discussion in specific empirical findings and discuss what these results imply for developers, practitioners, and governance of LLM-based multi-agent systems.
5.1. Personas Constrain Strategic Behavior and Bias Equilibrium Selection
As shown in Section 4.1, personas prevent models from achieving payoff-optimal equilibria in Tragedy-dominant scenarios, with equilibrium rates remaining near zero across models (Fig. 2) despite explicit payoff information. At the same time, as shown in Section 4.2, personas bias equilibrium selection toward Green Transition outcomes: with personas present, nearly all Nash equilibria correspond to Green Transition (Fig. 3), even in Tragedy-dominant scenarios where Tragedy is payoff-optimal. Removing personas enables Qwen models to select Tragedy equilibrium, while Llama and Mistral remain locked into Green selection regardless of condition.
These findings indicate that personas are not cosmetic enhancements but design choices that fundamentally alter system behavior. They shape both whether equilibrium is reached and which equilibrium is selected, consistent with prior findings that role prompts systematically bias strategic decision-making in LLMs (Neumann et al., 2025a; Herr et al., 2024; Hu et al., 2025; Macmillan-Scott and Musolesi, 2024), and with observations that LLMs exhibit cooperative or norm-aligned behavior even when such actions are payoff-dominated (Akata et al., 2025; Fontana et al., 2025; Lorè and Heydari, 2024; Fish et al., 2025).
Whether this bias is harmful depends on the deployment objective. If practitioners want incentive-responsive agents, persona framing interferes. If they want value-aligned agents, persona framing may help. The critical issue is that persona assignment makes this choice without explicit awareness, and the resulting system behavior may not match the designer’s intent.
This creates a risk of misinterpretation: systems that appear to simulate rational strategic behavior may instead encode designer-imposed normative assumptions through persona choices. Equilibrium outcomes in LLM-based simulations should not be interpreted as neutral predictions of system dynamics, but as outputs shaped by embedded normative assumptions, a concern echoed in work on LLM-based social simulation and sociotechnical impacts of AI systems (Aher et al., 2023; Gao et al., 2024; Domínguez Hernández et al., 2024; Rakova and Dobbe, 2023).
5.2. Payoff-Aligned Behavior Requires Both Persona Removal and Explicit Payoffs